OpenStudy (anonymous):

a car was bought new for 15 000. each year the value of the car depreciates by 30% of its value of the previous year. in how many years will the car be worth only $500.?

OpenStudy (amistre64):

A = P(1-r)^t

OpenStudy (amistre64):

A/P = (1-r)^t log[1-r](A/P) = t ln(A/P) ----- = t ln(1-r) if i did it right

OpenStudy (chaise):

Is this not a geometric progressions?

OpenStudy (amistre64):

to compact it for google: (ln(500/15000))/ln(1-.3) = abt 9.54

OpenStudy (chaise):

\[a _{n}= an ^{n-1}\]

OpenStudy (amistre64):

more like regression maybe :)

OpenStudy (amistre64):

to dbl check 500 = 15000(1-.3)^(9.54) 500 = 499.26 is pretty good

OpenStudy (chaise):

You answer seems pretty nice. I don't know how you got it - Although I'm 90% sure this problem can be solved using geometric progression. I just can't remember how it is done.

OpenStudy (amistre64):

the A= P(1-r)^t is the derived from the progression, but that usually aint covered till later on in the maths

OpenStudy (amistre64):

i just algebrad it into submission :)

OpenStudy (anonymous):

it doesn't have natural logs or progression because i haVen't learned that yet

OpenStudy (amistre64):

a0 = 15000 a1 = 15000(1-.3) a2 = 15000(1-.3)(1-.3) a3 = 15000(.7)^3 a4 = 15000(.7)^4 an = 15000(.7)^n 500 = 15000(.7)^n , and solve for n

OpenStudy (amistre64):

you have to know logs in order to solve for an exponential variable

OpenStudy (amistre64):

the only other option is to brute math it till it succombs :)

OpenStudy (anonymous):

the answer is 9.5 years

OpenStudy (amistre64):

yay!! .. i was close :) actually thats an approximation, but yes

OpenStudy (anonymous):

yes